A car travels at a constant rate of 40 miles per hour for 45 minutes. How far did the car travel?
Answer by jim_thompson5910(35256)

45 minutes = 45/60 = 0.75 hours

D = r*t

D = 40*0.75

D = 30

The car has traveled 30 miles.